OpenStudy (anonymous):

What is the measure of a side of a square whose diagonal is 4 inches longer than it's side? Round your answer to the nearest tenth.

OpenStudy (anonymous):

Let the side be x inches Therefore x^2 + x^2 = ( x + 4) ^2 2x^2 = x^2 + 8x + 16 x^2 - 8 x -16 =0 ( x-4) ( x-4) =0 x = 4 inches